August 2011 Reports

To the main

31st August 2011 (Wednesday)

Experiment activities:

  • Processing and stripping is finished. Nothing to report.

New GGUS (or RT) tickets:

  • T0: 0
  • T1: 3
  • T2: 0

Issues at the sites and services

30th August 2011 (Tuesday)

Experiment activities:

  • Processing and stripping is finished. Nothing to report.

New GGUS (or RT) tickets:

  • T0: 0
  • T1: 1
  • T2: 0

Issues at the sites and services

  • T0 *

29th August 2011 (Monday)

Experiment activities:

  • Processing and stripping is finished. Nothing to report.

New GGUS (or RT) tickets:

  • T0: 0
  • T1: 1
  • T2: 0

Issues at the sites and services

  • T0
  • T1

26th August 2011 (Friday)

Experiment activities:

  • Ongoing processing of data. finalisation of current production to clear old problem of data access

New GGUS (or RT) tickets:

  • T0: 0
  • T1: 1
  • T2: 0

Issues at the sites and services

25th August 2011 (Thursday)

Experiment activities:

  • Ongoing processing of data. finalisation of current production to clear old problem of data access

New GGUS (or RT) tickets:

  • T0: 0
  • T1: 1
  • T2: 0

Issues at the sites and services

  • T0
  • T1
    • CERN : file outside SPACE token (GGUS:73810)
    • SARA : is it possible to assigne the 1TB of "historical" LHCB-USER space token to LHCB-USER (GGUS:73087) we did not recived any INFO about the UNSCHEDULED downtime of SRM ....

24th August 2011 (Wednesday)

Experiment activities:

  • Ongoing processing of data. finalisation of current production to clear old problem of data access

New GGUS (or RT) tickets:

  • T0: 0
  • T1: 0
  • T2: 0

Issues at the sites and services

  • T0
  • T1

23rd August 2011 (Tuesday)

Experiment activities:

  • Ongoing processing of data.

New GGUS (or RT) tickets:

  • T0: 0
  • T1: 0
  • T2: 1

Issues at the sites and services

  • T0
  • T1
    • PIC :
    • SARA :
      • Problem of space token for lhcb-user 5 space tokens with 2 with DATA but only 1 active) (GGUS:73087)

22 August 2011 (Monday)

Experiment activities:

  • Ongoing processing of data.

New GGUS (or RT) tickets:

  • T0: 0
  • T1: 0
  • T2: 1

Issues at the sites and services

  • T0
  • T1
    • IN2P3 :
      • Problem "slow access to data" solved
    • PIC :
      • Problems with data access.
      • "Problems transferring files to CERN" (GGUS ticket 73576) was solved.
    • GRIDKA :
      • Faulty connections from 192.108.46.248 (GGUS 73630) still opened.

19 August 2011 (Friday)

Experiment activities:

  • Ongoing processing of data. Waiting for new data.

New GGUS (or RT) tickets:

Issues at the sites and services

  • T0
  • T1
    • Continuing problems at IN2P3 with access to data. Overloaded pool. GGUS ticket opened as requested.
    • PIC :
      • Continuing problems with PIC - degraded tape system. Many ongoing problems with access ti data there.
      • Problems transferring some files to CERN (GGUS ticket 73576). This is blocking further transfers from PIC to CERN.
    • GridKa :
      • Abnormally terminated connections from GridKa IP 192.108.46.248 overloaded some services in DIRAC this morning. These services have been restored, but there is the danger that other services can be affected in future. GGUS ticket opened.
      • Also problems with pilots aborted at cream-5-kit.gridka.de - GGUS ticket opened.

18 August 2011 (Thursday)

Experiment activities:

  • Ongoing processing of data

New GGUS (or RT) tickets:

  • T0: 0
  • T1: 0
  • T2: 2 (#73589 RALPP, #73590 Oxford)

Issues at the sites and services

  • T0
  • T1
    • Continuing problems at IN2P3.
    • Continuing problems with PIC - degraded tape system. Any update?
    • Somme user jobs failing access at GridKa - wait and see.
    • Accessing data at NIKHEF and SARA fine now.
    • GGUS problem sorted last evening.

17 August 2011 (Wednesday)

Experiment activities:

  • Ongoing processing of data

New GGUS (or RT) tickets:

  • T0: 0
  • T1: 0
  • T2: 1 (Aborted pilots at INFN CAGLIARI #73582)

Issues at the sites and services

  • T0
  • T1
    • Continuing problems at IN2P3 (more power cuts?)
    • Continuing problems with PIC - problem with their tape system. Any update?
    • Possible problems accessing data at NIKHEF and possibly SARA this morning. Wait and see.

16 August 2011 (Tuesday)

Experiment activities:

New GGUS (or RT) tickets:

  • T0: 0
  • T1: 1 (SARA #73478)
  • T2: 0

Issues at the sites and services

  • T0
  • T1
    • IN2P3 power cut problem affected many jobs. Seems to be recovering now.
    • Continuing problems with PIC - problem with their tape system.
    • SARA has added some capacity to LHCb-Disk. Looking forward to rest of disk coming online.
    • 3 Files lost at RAL due to bad tape. Recovery ongoing by LHCb.

15 August 2011 (Monday)

Experiment activities:

  • Quiet weekend. Backlog of jobs at GRIDKA cleared now. Running out of disk space (LHCb-Disk) at SARA. High number of problems with "input data resolution" at PIC.

New GGUS (or RT) tickets:

  • T0: 0
  • T1: 0
  • T2: 0

Issues at the sites and services

  • T0
    • AFS problems this morning
  • T1

12 August 2011 (Friday)

Experiment activities:

  • Data processing and stripping. We still have backlog of jobs at GRIDKA.

New GGUS (or RT) tickets:

  • T0: 0
  • T1: 0
  • T2: 0

Issues at the sites and services

  • T0
    • The CASTOR glitch reported yesterday (INC:058607) was understood: due to some incorrectly finished operation we had few files with zero size at CASTOR. We tried infinitely to send these files to CERN and all these transfers failed. After removing these corrupted files we have no any problems.
  • T1

11 August 2011 (Thursday)

Experiment activities:

  • Data processing and stripping. Backlog of waiting jobs at GRIDKA.

New GGUS (or RT) tickets:

  • T0: 0
  • T1: 0
  • T2: 0

Issues at the sites and services

10 August 2011 (Wednesday)

Experiment activities:

  • Data processing and stripping. Backlog of waiting jobs at GRIDKA.

New GGUS (or RT) tickets:

  • T0: 0
  • T1: 1
  • T2: 0

Issues at the sites and services

  • T0
  • T1
    • GRIDKA: Read only Shared Area for SAM jobs with software installation (GGUS:73343); Fixed

9 August 2011 (Tuesday)

Experiment activities:

  • A lot of data last day.
It was discovered RAWIntegrityAgent stuck, as result data transfer from pit to Castor was blocked. Fixed by restarting agent. FTS transfer from CERN to GRIDKA was blocked starting from weekend due to few wrong requests. No problems after requests were removed. Backlog of not processed files at GRIDKA.

New GGUS (or RT) tickets:

  • T0: 0
  • T1: 0
  • T2: 0

Issues at the sites and services

8 August 2011 (Monday)

Experiment activities:

New GGUS (or RT) tickets:

  • T0: 0
  • T1: 2
  • T2: 0
Issues at the sites and services

5 August 2011 (Friday)

Experiment activities:

New GGUS (or RT) tickets:

  • T0: 0
  • T1: 0
  • T2: 0
Issues at the sites and services
  • T0
    • Problems re-appearing for setting up runtime environment on batch nodes. This seems to be specific to a certain type of worker-node only (GGUS:73177)
    • Castor SRM down tonight and fixed swiftly this morning (GGUS:73213)
  • T1
    • SARA: several files found to be outside the space token, list of files will be provided by site (GGUS:73087). Another GGUS:73196 where a user cannot remove his files is probably related to the issue

4 August 2011 (Thursday)

Experiment activities:

New GGUS (or RT) tickets:

  • T0: 0
  • T1: 0
  • T2: 0
Issues at the sites and services
  • T0
    • Problems re-appearing for setting up runtime environment on batch nodes. This seems to be specific to a certain type of worker-node only (GGUS:73177)
  • T1
    • RAL: few files missing on the storage element. Currently under investigation by the site
    • SARA: several files found to be outside the space token, currently checking how to resolve the issue (GGUS:73087)

3 August 2011 (Wednesday)

Experiment activities:

New GGUS (or RT) tickets:

  • T0: 0
  • T1: 0
  • T2: 0
Issues at the sites and services
  • T0
  • T1
    • RAL: few files missing on the storage element. Currently under investigation by the site
    • SARA: several files found to be outside the space token, currently checking how to resolve the issue (GGUS:73087)

2 August 2011 (Tuesday)

Experiment activities:

New GGUS (or RT) tickets:

  • T0: 0
  • T1: 0
  • T2: 0
Issues at the sites and services
  • T0
    • Castor: Pending transfers to LHCb-Archive ST token are progressing, Current status is 10 files left to be transferred (INC:055007)
  • T1
    • SARA: Downtime this morning passed unnoticed
    • SARA: Thanks for quick handling of disk space allocation (GGUS:73090)

1 August 2011 (Monday)

Experiment activities:

New GGUS (or RT) tickets:

  • T0: 0
  • T1: 0
  • T2: 0
Issues at the sites and services
  • T0
  • T1 *

-- RobertoSantinel - 01-Jul-2011

Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r2 - 2011-09-01 - JoelClosier
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LHCb All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback